OpenStudy (anonymous):

I need help this is sooo hard 8th grade math!

OpenStudy (anonymous):

it is hard, isn't it?

OpenStudy (anonymous):

aalright haha im sure if you tae it in steps everything willl be simplar

OpenStudy (anonymous):

What is the value of the expression \[ \frac{ 1 }{ 4^{-3}}\]

OpenStudy (anonymous):

alright easy lets start by looking at the problem

OpenStudy (anonymous):

a.12 b.64 c. 1/64 d. 1/12

OpenStudy (anonymous):

we have a 4^-3 in the denominator. what does that tell you?

OpenStudy (anonymous):

Its either a or b

OpenStudy (anonymous):

umm idk

OpenStudy (anonymous):

lets not worry about the answer right now lets worry about the problem itself

OpenStudy (anonymous):

alright i will let you know a nice rule. when you bring a number from the denominator to the numerator you slip the exponent

OpenStudy (anonymous):

so something like 1/5^-2 = 5^2

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok thats fine. its just a rule. i just want to make sure you know 5^2 is 5 squared

OpenStudy (anonymous):

do you understand that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright cool

OpenStudy (anonymous):

so with your problem we have 1/4^-3

OpenStudy (anonymous):

we bring the 4^-3 to the numerator

OpenStudy (anonymous):

and we get?....what?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

...

OpenStudy (anonymous):

ok with our rule we change the sign of the exponent

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

4* -3 ???????

OpenStudy (anonymous):

oh i see what you are confused about. when i am talking about sign i mean positive or negative

OpenStudy (anonymous):

so all i mean by change the sign of the exponent is change it from -3 to 3

OpenStudy (anonymous):

so you have just 4^3

OpenStudy (anonymous):

now what is 4 to the power of 3?

OpenStudy (anonymous):

Its 64!

OpenStudy (anonymous):

tysm for helping me!

OpenStudy (anonymous):

you are welcome please look over notes on this

OpenStudy (anonymous):

its essential on later math problems you will have i highschool

OpenStudy (anonymous):

SO happy!

OpenStudy (anonymous):

im in the us

OpenStudy (anonymous):

well i mean like algebra and pre calc it makes things easier if you know basic stuff like this